### v2

# FastLGS: Speeding up Language Embedded Gaussians with Feature Grid Mapping

Abstract.
The semantically interactive radiance field has always been an appealing task for its potential to facilitate user-friendly and automated real-world 3D scene understanding applications. However, it is a challenging task to achieve high quality, efficiency and zero-shot ability at the same time with semantics in radiance fields. In this work, we present FastLGS, an approach that supports real-time open-vocabulary query within 3D Gaussian Splatting (3DGS) under high resolution. We propose the semantic feature grid to save multi-view CLIP features which are extracted based on Segment Anything Model (SAM) masks, and map the grids to low dimensional features for semantic field training through 3DGS. Once trained, we can restore pixel-aligned CLIP embeddings through feature grids from rendered features for open-vocabulary queries. Comparisons with other state-of-the-art methods prove that FastLGS can achieve the first place performance concerning both speed and accuracy, where FastLGS is 98 faster than LERF and 4 faster than LangSplat. Meanwhile, experiments show that FastLGS is adaptive and compatible with many downstream tasks, such as 3D segmentation and 3D object inpainting, which can be easily applied to other 3D manipulation systems. The project page is at https://george-attano.github.io/FastLGS/.
Teaser.
1. Introduction
With the potential in fields such as robotics and augmented reality, 3D scene understanding has always been a conspicuous topic in the research community. Given a set of posed images, the goal is to learn an effective and efficient 3D semantic representation along with scene reconstruction, which should naturally support a wide range of downstream tasks in scene manipulation. However, it remains a vital challenge to simultaneously achieve accuracy, efficiency and the grounding of open-vocabulary level semantics together.
Predominantly, existing methods usually only focus on the accuracy of differentiating objects, which normally assign points with labels or low-level features. For example, Semantic-NeRF (Zhi et al., 2021), N3F (Tschernezki et al., 2022) and Panoptic Lifting (Siddiqui et al., 2023) successfully lift labels or noisy 2D features into 3D radiance fields, providing semantic segmentations in certain scenes. However, their labels and low-level features (e.g., DINO (Caron et al., 2021) features) are not sufficient for natural modality interactions since some methods require extra position information (e.g., N3F(Tschernezki et al., 2022)) and others even do not provide any interaction (e.g., Semantic-NeRF(Zhi et al., 2021) and Panoptic Lifting(Siddiqui et al., 2023)), and they also suffer from poor quality concerning in-the-wild scenes due to the lack of zero-shot ability. One possible solution for introducing zero-shot semantics with friendly interactions is to use pre-trained text-image models with high-level features. LERF (Kerr et al., 2023), as one of the state-of-the-art examples, enables pixel-aligned queries of the distilled 3D CLIP (Radford et al., 2021) embeddings, supporting long-tail open-vocabulary queries. Nevertheless, despite solving the problem of open-vocabulary interactions, LERF’s results are completely not object-centric (in particular, it has been difficult to selectively edit specific regions or objects(Kobayashi et al., 2022)) and can only extract fuzzy relevancy maps, which might degrade in novel views because its non-object-centric CLIP ground truth under interpolation is neither 3D consistent nor accurate. The above discussion shows the dilemma of balancing open-vocabulary ability and localizing accuracy when building a semantic field, not to mention these NeRF-based methods also suffer from the slow rendering speed.
To get out of such a predicament, one may naturally think of achieving object-centric along with CLIP embeddings, but this brings numerous problems. First, we need to achieve the extraction of objects across all views to ensure object level 3D-consistency, where the popular Segment Anything Model (SAM) (Kirillov et al., 2023) is strong enough to serve. However, to be a zero-shot method without any other model guidance in the whole process, the raw multi-granularity SAM masks can be confusing for the many-to-many masks mapping of even a single object. Recent LangSplat (Qin et al., 2024) trains an auto-encoder to handle this issue, but the reconstructed features are degraded and not sufficient for demanding query tasks. Meanwhile, directly building a high-level CLIP field is not only time-consuming but also inconsistent across views because different views of one object can provide close but distinct CLIP features. Moreover, the query strategy is also largely determined by the building of a semantic field. If we query by computing the CLIP relevancy of each pixel as in LERF, it is still not affordable. These are all key challenges for building the interactive 3D semantic field.
Accordingly, in this paper, we present the FastLGS, which speeds up to build accurate 3D open-vocabulary semantic fields within 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) via the feature grid mapping strategy. Given a set of posed images, we ensure object-centric by using a mask-based method as the basis. FastLGS first extracts all SAM masks and sent through the CLIP encoder to extract object-level image features. While directly training a CLIP feature field in high dimensionality could be time-consuming, we build a feature grid to map CLIP features to lower 3D space, but which inherently creates a many-to-one mapping problem. To tackle the problem of many-to-one mapping problem in 3D-consistent mapping, we propose a cross-view feature matching strategy to assign feature grids that represent scene objects with all sides of semantics. The low-dimensional mappings are attached to Gaussians to build a feature field. During the inference stage, pixel-aligned features will be rendered and we restore their semantics within feature grids to generate results upon open-vocabulary queries.
During experiments, we found FastLGS can generate competitive target masks compared with state-of-the-art 3D segmentation and semantic field methods, grounding much more accurate language embeddings than LangSplat (see Figure 2). Meanwhile, the FastLGS is fast and supports real-time interactions under high resolution (98 faster than LERF and 4 faster than LangSplat, see Table 4), which can also be queried for single or multiple objects with similar queries in an adjustable way. Notably, FastLGS is adaptive and compatible with many downstream tasks that can be easily applied to systems like 3D segmentation and manipulation.
In summary, the principal contributions of this work include:
To our best knowledge, this is the first work that can interactively generate scene target masks upon natural language queries under high resolution (14401080) in real-time.
We propose the grid mapping strategy with the zero-shot open-vocabulary prompted paradigm for building a 3D semantic field in 3DGS. As far as we know, our method is one of the fastest and most accurate methods in generating pixel-aligned semantic features in the 3D space.
We provide an efficient 3D semantic basis and find it adaptive to easily integrate with several downstream tasks like 3D segmentation and 3D object inpainting.
2. Related works
2.1. NeRF and 3D Gaussian Splatting (3DGS)
Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3DGS) have revolutionized 3D scene modeling and rendering. Introduced by (Mildenhall et al., 2021), NeRF synthesizes photorealistic scenes from sparse images by modeling volumetric density and color, further enhanced by innovations like mip-nerf (Barron et al., 2021) and TensoRF (Chen et al., 2022). 3D Gaussian Splatting, detailed by (Kerbl et al., 2023) and expanded upon in work (Luiten et al., 2024) on dynamic 3D Gaussians, optimizes the rendering of point clouds with Gaussian kernels for real-time applications. These developments continue to advance the field, highlighting the dynamic growth and application breadth of these technologies. Additional contributions by (Tang et al., 2023b) and others (Fang et al., 2022) underscore the ongoing enhancements and versatility of 3DGS in handling increasingly complex rendering tasks. Noting the advantages of 3D Gaussian Splatting in rendering speed and quality, our paper chooses it as a representation for building semantic field.
2.2. 2D and 3D Segmentation
The field of 2D image segmentation has undergone remarkable advancements since the introduction of Fully Convolutional Networks (FCN) (Long et al., 2015), which transformed traditional Convolutional Neural Networks (CNN) into architectures capable of pixel-wise predictions, revolutionizing segmentation tasks. This innovation spurred expansive growth, with significant contributions like Mask R-CNN (He et al., 2017), which added a branch for object mask prediction to Faster R-CNN, and the adoption of Transformer architecture, notably through SEgmentation TRansformer (SETR) (Zheng et al., 2021), alongside studies (Cheng et al., 2021; Strudel et al., 2021; Xie et al., 2021; Cheng et al., 2022). These works demonstrated the Transformer’s potential in capturing complex dependencies for enhanced semantic segmentation. Recent innovations such as SAM (Kirillov et al., 2023) and SEEM (Zou et al., 2023) utilize various kinds of prompts for segmentation, inspired many mask based researches.
Advancements in 3D segmentation have paralleled those in 2D, with significant breakthroughs such as PointNet (Qi et al., 2017a) and PointNet++ (Qi et al., 2017b), which revolutionized the direct processing of point sets for classification and segmentation. These technologies enhanced the handling of 3D data by adapting and scaling to capture intricate local structures. Further innovations include Recurrent Slice Networks (RSNets) (Huang et al., 2018) for improved volumetric data segmentation, and Cylinder3D (Zhou et al., 2020), tailored for LiDAR semantic segmentation in driving scenes. Following developments in radiance fields have also refined the precision and applicability of 3D segmentation techniques (Goel et al., 2023; Tang et al., 2023a; Cen et al., 2023b), but open-vocabulary text guided 3D segmentation remains a challenge for the lack of compatible semantic scene construction.
2.3. Semantic 3D scene
Advancements in semantic integration have significantly enhanced 3D modeling, notably through NeRF and 3DGS. (Zhi et al., 2021) introduced semantic layers into NeRF, setting the stage for enriched scene understanding. This has been further developed in studies such as LERF (Kerr et al., 2023), ISRF (Goel et al., 2023), DFF (Yen-Chen et al., 2022), and N3F (Tschernezki et al., 2022), which integrate detailed semantic segmentation to refine 3D visualizations.
In parallel, enhancements in 3DGS have also advanced semantic capabilities. The foundational work (Kerbl et al., 2023) and further developments (Gu et al., 2024; Ye et al., 2023) have refined the application of 3DGS in semantic segmentation tasks. The SAGA framework (Cen et al., 2023a) demonstrates the integration of detailed 2D segmentation outcomes into 3D models, enhancing semantic accuracy. LangSplat (Qin et al., 2024) explores incorporating language processing into 3DGS, adding a new layer to semantic interpretations in 3D scene modeling. However, with inconsistent or fuzzy results in varying views and slow query speed, the quality of semantic fields built by these approaches still limits their further applications.
Pipeline of FastLGS
3. Method
3.1. Overview
As in Figure 3, the input images are initially put for extracting SAM masks and CLIP embeddings (Sec 3.2). This information is utilized together for the building of a semantic feature grid and low-dim feature mapping (Sec 3.3 and Sec 3.4). Then, the obtained pixel-aligned low-dim features are learned in parallel with scene reconstruction according to Sec 3.5 where we describe the training for embedded features through Gaussians. At the inference stage, the built semantic field can be interactively queried for high quality masks in novel views using natural language based on restored embeddings. The query strategy is demonstrated in Sec 3.6.
3.2. Initialization
First, we initialize the original semantics. Given input images , for each image , we utilize SAM to obtain segmentation masks with a regular grid of 32 32 point prompts and compute respective CLIP features . Masks and CLIP embeddings will be used for building feature grids along with pixel-aligned low-dim feature maps. After initialization, no other pre-trained model will be required except for a CLIP text encoder at the inference stage.
3.3. Semantic Feature Grid
By practice, simply integrating CLIP feature with 3DGS can lead to infeasible demand of memory and limited efficiency of rendering as each Gaussian needs to record and update a parameter in hundreds of dimensions (Sec 4.3). On the other hand, the CLIP features of an object could show unexpected variance due to the view angles, redundant background and complex occlusion, posing a challenge to maintaining consistency and accuracy of semantics in 3D scenes. As language features of the same object from multiple views should share similar semantics, it could be compressed for lower training and query costs. However, current MLP-based compressed and reconstructed features cannot provide queries of higher quality as the former can be inconsistent and inaccurate like the raw features while the latter can be degraded which is not sufficient for demanding scenes (see relevancy maps in Figure 2 and Figure 6). Meanwhile, similar methods also can not acquire unseen features in a single view, restricting the possibility of querying objects with semantics captured in another view but invisible through the current angle.
To address these issues, we propose semantic feature grid, which stores the multi-view language features of each object in the scene and is mapped to a low dimensional feature. Specifically, our goal is to map to where language features L of the same object share the same low-dim feature f. Each feature f is assigned to a grid with the size of , where refers to the number of objects in a specific scene computed during cross view matching (Sec 3.4). The low dimensional features will be assigned pixel-wise according to corresponding masks and trained efficiently in 3DGS, while the grid can be used to restore accurate CLIP embeddings based on rendered low-dim features during inference. In practice, this strategy enables our method to achieve less consumption and faster speed (Table 4).
3.4. Cross View Grid Mapping
To achieve consistent features of objects across views, we sequentially match the denoised masks in adjacent image set I and assign features. Considering the possible instability of CLIP features in multi-views mentioned in Sec 3.3, the matching process consists of both key points correspondence and feature similarity comparison.
Key points correspondence We prioritize the mask correspondence by key points obtained using the SIFT and the k-nearest neighbors (KNN) algorithm for accurate and robust matching. The SAM masks with more than corresponding key point pairs will be assigned to the same low dimensional feature f and considered as segmentation masks of an object from different viewpoints.
Feature similarity comparison As some segmentations with smooth pixel value have few key points, the matching process is employed according to similarity calculated by hybrid features of CLIP embedding L and color distribution C. For and with and mismatched masks, we calculate the similarity matrix and choose the mask pairs with highest similarity. A threshold is set that masks with similarity higher than will have the same feature f and the others will be assigned to a new low dimensional feature.
For two segmentation masks , , the similarity is calculated as follow:
where is the weight of similarity of color distribution features C. is calculated based on Bhattacharyya distance of C while is calculated by cosine similarity:
Algorithm 1 displays our procedure of correspondence-based grid feature mapping. The consistency of objects across views is built and the amount is accessed during the matching process, which helps establish our semantic feature grids and low-dimensional feature mapping.
3.5. Training Features for Gaussians
We train a mapped low-dimensional feature for each gaussian to build a feature field, which is adaptive for integration.
Rendering Features Given a camera pose , we compute the feature of a pixel by blending a set of ordered Gaussians overlapping the pixel similar to the color computation of 3DGS:
where is given by evaluating a 2D Gaussian with covariance multiplied with a learned per-Gaussian opacity.
Optimization The low-dim features follows 3DGS optimization pipeline and especially inherits the fast rasterization for efficient optimization and rendering. The loss function for features is combined with a D-SSIM term:
where is also fixed to 0.2 in all cases.
3.6. Querying Features
Once trained, FastLGS can be interactively queried for different objects in the scene using open vocabulary text prompts. First, we query the pixel-aligned semantics by projecting the attached feature to view the plane using the rendering for features mentioned in Sec. 3.5, which can be used for all queries within camera pose .
Relevancy Score When provided a prompt, we assign its language relevancy scores based on grid image features similar to LERF. We compute the cosine similarity between image embedding and canonical phrase embeddings , then compute the pairwise softmax between image embedding and text prompt embedding , so that the relevancy score is:
For canonical phrases, we use “object”, “stuff” and “texture” for all queries.
Target Mask We go through feature grids to restore high-level image features and use them to compute relevancy with text embeddings. The grid with the highest relevancy score is deemed to be the target grid queried. Because each grid has multi-view image features of an object, we are even able to locate targets based on semantics invisible in the current query view (for example, locating a book named “computer vision” from the back of it). While the rendered semantics have been smoothed, we locate their corresponding feature grids by calculating euclidean distance between grids and rendered features: where denotes the grid’s corresponding low-dim features. We compute pixel-wise based on target grid, and points with lower than threshold forms the target mask. By changing the number of grids used, we can adjustably query multiple targets (as Figure 1).
4. Experiments
In this section, we first show the speed and quality of open-vocabulary object retrieval in comparison with other state-of-the-art methods through quantitative experiments in Sec 4.1, then we provide qualitative results of downstream tasks including 3D segmentation and object deletion in Sec 4.2. Ablation studies are in Sec 4.3 to demonstrate the rationality of feature grid based design.
Datasets For quantitative experiments, we train and evaluate the models on datasets including SPIn-NeRF (Mirzaei et al., 2023), LERF (Kerr et al., 2023) and 3D-OVS (Liu et al., 2023). For experiments on downstream applications we also used the MIP-360 datatset (Barron et al., 2022) (see example 4 of Figure 6).
Implementation Details We use the same OpenClip ViT-B/16 model as LERF and SAM ViT-H model as LangSplat. We train the features and scenes in 3DGS for 30,000 iterations. Activation threshold , correspondence threshold and are set to , 4 and 0.95. Weight is set to 0.3. f is normalized to . Experiments show robustness of the proposed method to the aforementioned parameters. While the original time calculation of LangSplat puts aside feature rendering time and feature reconstructing time, here we compute the query time of the whole query process as LERF does. The tested LERF masks are regions with relevancy higher than 20% after normalization. The normalization for each query is from 50% (less relevant than canonical phrases) to the maximum relevancy, which is identical to the visualization strategy of LERF. All results are reported running on a single TITAN RTX GPU.
Qualitative visual results on different datasets.
4.1. Comparison With Other State-of-the-art Methods
For quantitative experiments, we test different methods on the SPIn-NeRF dataset, the LERF dataset and 3D-OVS dataset. We show the quality of FastLGS-generated masks by comparing with other state-of-the-art 3D segmentation methods including the multi-view segmentation of SPIn-NeRF (MVSeg) (Mirzaei et al., 2023) and SA3D (Cen et al., 2023b), and compare the language retrieval ability with LERF (Kerr et al., 2023) and LangSplat (Qin et al., 2024).
SPIn-NeRF dataset We evaluate the IoU and pixel-wise accuracy of masks with provided ground truth (1008 567), and also show the time consumption for each query, which is omitted in MVSeg and SA3D because they do not support single view query. For LERF, LangSplat and FastLGS, we use the same text queries for the target segmentation objects to generate masks. Other methods all follow their original settings when tested on this dataset. We also provide the 2D segmentation results generated by SAM based on manual point prompts of original scene images for comparison. As shown in Table 1, FastLGS generates masks with competitive quality which have higher mIoU compared with other methods, and largely outperforms LERF and LangSplat in query speed.
LERF dataset The LERF dataset contains a number of in-the-wild scenes and is much more challenging, which strongly requires zero-shot abilities. We report localization accuracy for the 3D object localization task following LERF (Kerr et al., 2023) with ground truth annotations provided by LangSplat (Qin et al., 2024) (resolution around 985 725). Results are shown in Table 2 and visual examples in Figure 4 (line 1&2), which further demonstrates FastLGS’s advantages in natural language retrieval.
3D-OVS dataset We also compare with 2D-based open-vocabulary segmentation methods including ODISE (Xu et al., 2023) and OV-Seg (Liang et al., 2023) along with 3D-based methods including 3D-OVS (Liu et al., 2023), LERF (Kerr et al., 2023) and LangSplat (Qin et al., 2024). Results are provided in Table 3, where our method can outperform both 2D and 3D methods. In Figure 4 (line 3) our method can also better support demanding custom query “keyboard and pocket tissues” which fails in LangSplat. Meanwhile, Figure 6 provides a visualization of pixel-aligned relevancy upon different queries, where our method has more accurate relevancy scores compared with LERF and LangSplat. Table 4 further shows the query time and model size where our method also outperforms LERF and LangSplat, proving its advantages in real-world applications.
| Method | mIoU (%) | mAcc (%) | mTime (s) |
| SAM(2D) | 95.7 | 96.1 | 0.05 |
| MVSeg | 89.5 | 97.8 | - |
| SA3D | 90.9 | 98.3 | - |
| LERF | 81.0 | 99.5 | 30.2 |
| LangSplat | 92.2 | 94.7 | 1.43 |
| LEGaussians | 89.3 | 97.8 | 1.04 |
| OURS | 93.1 | 95.2 | 0.31 |
| Scene | LERF | LangSplat | LEGaussians | OURS |
| ramen | 61.9 | 73.2 | 78.6 | 84.2 |
| figurines | 75.5 | 80.4 | 83.7 | 91.4 |
| teatime | 84.8 | 88.1 | 93.6 | 95.0 |
| waldo_kitchen | 70.2 | 95.5 | 95.1 | 96.2 |
| Overall | 73.1 | 84.3 | 87.8 | 91.7 |
| Method | bed | bench | room | sofa | lawn | Overall |
| ODISE | 55.6 | 30.1 | 53.5 | 49.3 | 39.1 | 45.5 |
| OV-Seg | 79.8 | 88.9 | 71.4 | 66.1 | 81.2 | 77.5 |
| 3D-OVS | 89.5 | 89.3 | 92.8 | 74.1 | 88.2 | 86.8 |
| LERF | 76.2 | 59.1 | 56.4 | 37.6 | 78.2 | 61.5 |
| LangSplat | 92.6 | 93.2 | 94.1 | 89.3 | 94.5 | 92.7 |
| LEGaussians | 45.7 | 47.4 | 44.7 | 48.2 | 49.7 | 47.14 |
| OURS | 94.7 | 95.1 | 95.3 | 90.6 | 96.2 | 94.4 |
| Scene “sofa” | LERF | LangSplat | OURS |
| mTime (s/q) | 51.2 | 2.14 | 0.52 |
| Model Size | 1.28GB | 565MB | 200MB |
4.2. Downstream Applications
In this section, we test our method for the ability of applying to downstream 3D object manipulation tasks. Successful visual results are provided in Figure 6.
Language Driven 3D Segmentation We integrate FastLGS with Segment Any 3D Gaussians (SAGA) (Cen et al., 2023a) to achieve language driven 3D segmentation. SAGA attaches a low-dimensional 3D feature to each Gaussian in the trained 3DGS model to conduct mask-based segmentation. We attach our low-dim features to Gaussians in parallel with SAGA. The original SAGA requires manual positional prompting for reference mask selection, but after integration we can directly use our masks as segmentation references and conduct 3D segmentation based on natural language. The whole query and segmentation process can be completed around only 1s.
Language Driven Object Inpainting We generate target masks of queried objects across views and use them as multi-view segmentation masks. We follow the multi-view segmentation based inpainting pipeline of SPIn-NeRF (Mirzaei et al., 2023) to conduct 3D object inpainting using the extracted multi-view segmentation masks. The masks are dilated by default with a 5x5 kernel for 5 iterations to ensure that all of the object is masked as in SPIn-NeRF. Results show that consistent masks can be directly used for high-quality object inpainting through 3D inpainting methods.
4.3. Ablations
We conduct ablation to validate the necessity of our semantic feature grid for query and the improvements for the accuracy of our components in building cross-view grid mapping.
Querying Feature The performance is reported in Table 5, where SFG represents the semantic feature grid. Initially, we directly reconstruct the semantic field with NeRF using extracted CLIP features for SAM-based masks of images. When replacing NeRF with 3D Gaussian Splatting, the high dimensions of CLIP feature result in running out of memory. The issue is addressed by combining it with a semantic feature grid, which also brings improvements in accuracy and speed.
Cross view We report results in Table 6 and Figure 5, where KP represents key points and CD represents color distribution feature C. Key points correspondence and features similarity with the auxiliary of color distribution provide more guidance for maintaining consistency of objects across a series of continuous perspectives, which are dedicated to training low-dim features of high quality for Gaussians and effectively improving the performance.
| Component | Performance | |||
| SAM | 3D-GS | SFG | mIoU(%) | mTime(s) |
| ✓ | 83.3 | 20.1 | ||
| ✓ | ✓ | OOM | OOM | |
| ✓ | ✓ | ✓ | 95.1 | 0.98 |
| Component | Performance | |||
| CLIP | KP | CD | mIoU(%) | mAcc(%) |
| ✓ | 75.2 | 83.2 | ||
| ✓ | ✓ | 87.5 | 91.8 | |
| ✓ | ✓ | ✓ | 92.1 | 97.5 |
This is ablation.
More visual results on different datasets. Example 1: Query relevancy visualization. Example 2&3: Object retrieval. Example 4: Downstream 3D segmentation. Example 5: Downstream 3D inpainting.
5. Conclusions and future work
In this paper, we propose FastLGS, a method that speeds up to build accurate 3D open-vocabulary semantic fields within 3DGS. By mapping and restoring multi-view CLIP embeddings through feature grids instead of MLPs, FastLGS not only has interactive query speed, but also provides more accurate target masks and semantic relevancy along with lower consumption compared with state-of-the-art methods. Meanwhile, experiments show FastLGS’s ability to be integrated with other downstream 3D manipulation tasks. While FastLGS now follows an object-centric paradigm, challenges occur in querying for parts of an object, such as “bicycle handlebars” or “chair leg”. We believe this could possibly be solved by building the feature grid in a multi-granularity way, where actual performance (time and space) along with cross-view consistency of both mask extraction and query results remain a problem, which requires further researches.